shift summary: owl 9/8/99

John Shields (jshields@ksera)
Wed, 8 Sep 1999 06:59:54 -0500

Shift Summary for owl shift, 09/08/99

On Shift: J. Shields, R. Kessler

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

tevatron down due to MI and sextupole power supply problems.
========================================================================
Runs ended this shift: NONE

Tapes brought to FCC this shift: NONE
(including tapes for online split)

========================================================================
Accesses during shift:

none
========================================================================
Overview of shift:

waiting for beam to return. And waiting, and waiting, and......
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

00:00 shift crew arrives and discovers that tevitron
has been down since 23:11 due to MI problems.
estimate was 4 hours.

01:00 RK fixes BEAM_832_REGOUT destination logic

03:21 Surprise! Beam doesn't come back as advertised.
Instead we are treated to a new chan 13 message:
"All power restored to both MI and Pbar. MI ramp
has been established. Beam is currently being held
off by a sextupole power supply problem. Experts
are en route.

03:30 RK switches the RF from Bob T's temporary RF back
to the normal tevatron RF.

14464 03:49 Shift crew very optimistically begins a new run
as a contingency against the FANTASICALLY unlikely
scenario that beam will actually return during
their shift.

06:20 A new channel 13 message pops up and dashes
the crew's feeble hope of seeing even a faint glimmer
of beam during their watch. Message text as follows:
"Vertical sextuple PS in MI holding off beam.
Repairs underway. Estimate is 3 hours."

========================================================================
Follow up Notes on Previous Shift

regular tevatron RF restored

========================================================================
Items to Follow up on Next Shift